OpenStudy (anonymous):

How would you calculate the integral of 3x/((x-a)(x-b))dx when a cannot be equal to b?

OpenStudy (freckles):

have you tried partial fractions

OpenStudy (anonymous):

I did partial fraction and got A = -3a/b-a and B = -3b/a-b, but when it comes to integrating everything together I get -3a(ln(x-a))/b-a -3b(ln(x-b))/a-b. My online homework says it's wrong but i think I did everything correctly. Unless the formatting if off?

OpenStudy (freckles):

\[\text{ your work looks fine } \\ A \ln |x-a|+B \ln|x-b|+C \text{ where } A=\frac{3a}{b-a} \text{ and } B=\frac{-3b}{a-b}\]

OpenStudy (freckles):

like when you put in did you put parenthesis around the denominators?

OpenStudy (freckles):

3a*ln|x-a|/(b-a)-3b*ln|x-b|/(a-b)

OpenStudy (anonymous):

Yes, I put parenthesis around the denominators, but even then I still keep getting it wrong.

OpenStudy (freckles):

We could try to write as one term also it might also help to out the | | around the thing that is in the ln

OpenStudy (anonymous):

I'll try changing the format of my answer and see if it works.

OpenStudy (freckles):

\[\frac{-3a}{a-b} \ln|x-a|+\frac{-3b}{a-b} \ln|x-b|+C \\ \frac{-3}{a-b}[a \ln |x-a|+b \ln|x-b|]+C \\ \frac{-3}{a-b}[ \ln|(x-a)^a|+\ln|(x-b)^b|+C \\ \frac{-3}{a-b}\ln|(x-a)^a(x-b)^b|+C\]

OpenStudy (freckles):

they it say anything special about how you should write it or simplify it

OpenStudy (freckles):

does it (not they)

OpenStudy (anonymous):

It only says to calculate for a when it isn't equal to b. I keep getting it marked incorrectly.

OpenStudy (freckles):

can you show me how you enter it

OpenStudy (freckles):

and are you doing ln( ) or ln| |

OpenStudy (anonymous):

It doesn't give me the option to do ln | |, it has to be ln ( ). My online homework is set up that way.

OpenStudy (anonymous):

3a(ln(x-a))/(b-a) - 3b(ln(x-a))/(a-b)

OpenStudy (freckles):

and you did put the +C

OpenStudy (anonymous):

yes

OpenStudy (freckles):

i notice when you wrote what you just wrote you put x-a in that second ln

OpenStudy (freckles):

if that isn't it i give up

OpenStudy (freckles):

should be 3a(ln(x-a))/(b-a) - 3b(ln(x-b))/(a-b)+C

OpenStudy (anonymous):

Oh. Typo. But yea I just plugged it in again and nothing. Thank you so much for the help though. I'll see how to get around this mess.

OpenStudy (freckles):

\[\frac{A}{x-a}+\frac{B}{x-b}=\frac{3x}{(x-a)(x-b)} \\ A(x-b)+B(x-a)=3x \\ (A+B)x+(-Ab-Ba)=3x \\ A+B=3 \text{ and } -Ab-Ba=0 \\ A=3-B \\ -(3-B)b-Ba=0 \\ -3b+Bb-Ba=0 \\ B(b-a)=3b \\ B=\frac{3b}{b-a}=\frac{-3b}{a-b} \\ A=3-B=3+\frac{3b}{a-b}=\frac{3(a-b)+3b}{a-b}=\frac{3a}{a-b}\]

OpenStudy (freckles):

maybe we wrote A wrong

OpenStudy (freckles):

I just got a different A

OpenStudy (anonymous):

Sorry for all the typos. I meant to put positive A.

OpenStudy (anonymous):

But when it comes to integrating 3a/a-b * 1/x-a, you move 3a/a-b to the front of the integral correct?

OpenStudy (freckles):

\[\frac{3a}{a-b} \ln|x-a|-\frac{3b}{a-b} \ln|x-b|+C\]

OpenStudy (anonymous):

It's still incorrect. Appreciate the help. I'm trying to see what the problem may be.

OpenStudy (freckles):

http://www.wolframalpha.com/input/?i=integrate%283x%2F%28%28x-a%29%28x-b%29%29%2Cx%29 this confirms my last answer there

OpenStudy (freckles):

i don't know what else could be the problem

OpenStudy (anonymous):

I'm not sure either. But thank you!

OpenStudy (freckles):

sorry couldn't be more help good luck